OpenStudy (anonymous):

A taxi ride cost $29.40. The driver charged $3 plus $0.40 per 0.2 mile traveled. How far did the taxi travel on this trip?

OpenStudy (anonymous):

0.2x * 0.40 = y 3x + y = 29.40 subtituting 3x + 0.2x * 0.40 = 29.40 3x+0.2x = 29.40 / 0.40 x = (29.40/0.40)/3.2

OpenStudy (anonymous):

$29.40-$3=$26.40 $26.40/0.4=66 66 x 0.2=13.2miles im pretty sure

OpenStudy (anonymous):

hmmmmm, i misread $3 plus $0.40 per 0.2 mile, garg u're right, sorry lisa

OpenStudy (anonymous):

Thanks. I makes sense now that I see it worked out.